There is a substance in a solution (8g/liter). The length of a cuvette is 1cm and only 25% of the certain light beam is transmitted. What is the absorption coefficient?
Answer #1

We have , Absorbance = log ( I 0 / I )

Where , I 0 intensity of incident light and I intensity of transmitted light.

We have, I 0 =100 % and I = 25 % .

Therefore, A = log (100 /25 ) = 0.602

From Beer's law , we have relation A = \epsilon b C

Where A is absorbance, \epsilon is absorption coefficient , b is path length and C is concentration of substance.

\therefore 0.602 = \epsilon\times 1 cm \times 8 g / L

\epsilon = 0.602 / 1 cm \times 8 g / L = 0.075 g L -1 cm -1

ANSWER : absorption coefficient of substance = 0.075 g L -1 cm -1